User name Higinbotham
Log entry time 10:29:50 on December 18,1999
Entry number 31393
keyword=E-HRS to 87.5 and 0.583 Mev
Electron Spectrometer settings changed to 87.4 degrees and magnets set
to 0.583 GeV/c. Since we are going up in momentum, we cycled the quads
by going to 1600 A, waiting 5 min., going to 1000 A, and then reset the momentum.